OpenStudy (anonymous):

I have to solve using elimination. x + y -2z = 8 5x - 3y + z = -6 -2x - y + 4z = -13 I think I'm doing it right, so far I've got -5(x + y -2z = 8) -5x - 5y + 10z = 8 5x - 3y + z = -6 _____________________ -8y + 11z = -46 from there, do I isolate y and then substitute the y into the other equation or?

OpenStudy (zzr0ck3r):

Have you heard of Gaussian elimination?

OpenStudy (anonymous):

No?

OpenStudy (zzr0ck3r):

I would multiply the first equation, by 2 and then add it to the last equation.

OpenStudy (zzr0ck3r):

then you get y=3

OpenStudy (anonymous):

So it would be 2(x + y -2z = 8) 2x + 2y -4z = 16 -2x - y + 4z = -13

OpenStudy (zzr0ck3r):

yeah

OpenStudy (anonymous):

so it cancels out both the x and z variables making it y=3?

OpenStudy (anonymous):

then I can substitute that into the the first equation right?

OpenStudy (zzr0ck3r):

yes, now plug that into the first two equations

OpenStudy (zzr0ck3r):

then you have x-2z=5 and 5x+z=3 can you solve this system of two equations with two variables?

OpenStudy (anonymous):

I think I can, let me try it out on paper if that's ok

OpenStudy (anonymous):

I did x + 3 -2z = 8 -2x - 3 + 4z = -13 ____________________ -x + 2z = -5 -x = -2z - 5 x = 2z + 5

OpenStudy (anonymous):

but then I did x + y -2z = 8 2z + 5 + 3 - 2z = 8 the z's cancel each other out so I think I did something wrong or it doesn't have a solution?

OpenStudy (zzr0ck3r):

no it does, if you take y=3 and plug into the first two equations, you get x-2z=5 and 5x+z=3 multiply the first one by -5 -5x+10z=-25 and we have the other one 5x+z=3 add them 11z=-22 z=-2

OpenStudy (zzr0ck3r):

now solve and get x=1, y=3, z=-2

OpenStudy (fibonaccichick666):

If you're interested, check out matrices-specifically the coefficient matrix- on khanacademy, you'll notice it is exactly what you are doing, but with less writing!

OpenStudy (zzr0ck3r):

yeah I mentioned Gaussian elimination at the top but if he didn't know about it I am guessing they should not use it. But it is much nicer...

OpenStudy (anonymous):

you kind of lost me when you got x-2z=5 and 5x+z=3 when I substituted, I got x + 3 - 2z = 8 and 5 -9 + z

OpenStudy (anonymous):

@FibonacciChick666 mentioned matrices, but we skipped that in the book

OpenStudy (fibonaccichick666):

they should really teach matrices instead of this crap. @fluttere, if you don't mind looking, gaussian elimination is worth learning

OpenStudy (fibonaccichick666):

@fluttere lol

OpenStudy (fibonaccichick666):

....really? It won't let me tag your name

OpenStudy (zzr0ck3r):

where I lost you plug in y=3 to the first equation and simplify x+y-2z=8 becomes x+3-2z=8 becomes x-2z=5 do the same thing for the second equation, plug in y=3 5x - 3y + z = -6 becomes 5x-9+z=-6 5x+z=3 so we have x-2z=5 5x+z=3

OpenStudy (zzr0ck3r):

now solve that, and get the two answers for x and z, and we know y=3

OpenStudy (zzr0ck3r):

Go read about matrix elimination method. It is the exact same thing in different notation, and you will be surprised how much the notation makes it easier to see.